The manufacturing capacity of Susil Company's facitilites is 30,000 untis of product a year. A summary or operating results for the year ended December 31, 2012 is as follows:

   Sales (18,000 units @ $100)                                    $1,800,000
   Variable manufacturing & selling costs ($55 per unit)    990,000
   Contribution Margin                                                 $ 810,000
   Fixed costs                                                                 495,000
   Operating Income                                                    $ 315,000

A foreign distributor has offered to buy 15,000 units at $90 per unit during 2013. Assume that all of Susil's costs would be at the same levels and rates in 2013 as in 2012. If Susil accepts this offer, it will have to reject some regular business from regular customers so as not to exceed capacity. If they acccept the special order, how much will operating income change?

Question 10 options:

a)

$525,000 increase

b)

$390,000 increase

c)

$675,000 decrease

d)

$705,000 increase
Answer #1

Present operating income will be = (15000*45)-495000 = 180000

Special order net income = (90-55)*15000 = 525000

Total Net income after special order = 525000+180000 = 705000

Previous operating income = 315000

Change in operating income = 705000-315000 = 390000

So answer is b) $390000 Increase